# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ A.2 Proofs of Transfer Learning
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f},h})]\leq 4\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]+C_{xy}\log^{3}(m)r_{x},$ (A.19)
@@
$\Phi=\{(x,y)\mapsto\widetilde{\ell}(x,y,f):=(\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}}))\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}:f\in\mathcal{F}\},$ (A.20)
where the truncation radius $R\geq 1$ will be defined later. It is easy to show that with probability no less than $1-2m\exp(-C_{1}^{\prime}R^{2})$, it holds that $\|x_{i}\|_{\infty}\leq R$ for all $1\leq i\leq m$. Hence by definition, the empirical minimizer also satisfies $\widehat{f}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)$. Below we reason conditioned on this event and verify the conditions required in Lemma A.11.
@@
    ~$\widetilde{\ell}(x,y,f)\leq\mathbb{E}_{t,x_{t}|x}\|s_{f,h}(x_{t},y,t)-\nabla_{x}\log\phi_{t}(x_{t}|x)\|^{2}\lesssim M_{f}^{2}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T-T_{0}}+1\Big{)}.$ (A.21)
    And by Lemma A.10,
    ~$-\widetilde{\ell}(x,y,f)\leq\mathbb{E}_{t,x_{t}|x}\|s_{*}^{\mathbb{P}}(x_{t},y,t)-\nabla_{x}\log\phi_{t}(x_{t}|x)\|^{2}\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}\lesssim C_{X}^{{}^{\prime\prime}}R^{6}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T}+1\Big{)}.$ (A.22)
    Let $M:=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T}+1\Big{)}\right)$ and thus |~$|\widetilde{\ell}(x,y,f)|\leq M$.
@@
    $\mathbb{E}_{(x,y)\sim\mathbb{P}}\left[\mathbbm{1}_{\|x\|_{\infty}\leq R}\left(\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})\right)^{2}\right]$ (A.23)
@@
    $\Big{\|}\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\sigma_{i}\widetilde{\ell}(x_{i},y_{i},f_{1})-\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\sigma_{i}\widetilde{\ell}(x_{i},y_{i},f_{2})\Big{\|}_{\psi_{2}}\leq 4\|\widetilde{\ell}(\cdot,\cdot,f_{1})-\widetilde{\ell}(\cdot,\cdot,f_{2})\|_{L^{2}(\widehat{\mathbb{P}}_{m})},$ (A.24)
    where ^$\widehat{\mathbb{P}}_{m}:=\frac{1}{m}\sum_{i=1}^{m}\delta_{(x_{i},y_{i})}$. Define $\Phi_{r}:=\{\varphi\in\Phi:\frac{1}{m}\sum_{i=1}^{m}\varphi(x_{i},y_{i})^{2}\leq r\}$ and it is easy to show that $\textbf{diam}\big{(}\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})}\big{)}\leq 2\sqrt{r}$. By Dudley’s bound [Van Handel, 2014; Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta>0$,
    $\mathcal{R}_{m}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})},\varepsilon)}{m}}\ \mathrm{d}\varepsilon\right).$ (A.25)
@@
    $\frac{1}{m}\sum_{i=1}^{m}(\widetilde{\ell}(x_{i},y_{i},f_{1})-\widetilde{\ell}(x_{i},y_{i},f_{2}))^{2}$ $=\frac{1}{m}\sum_{i=1}^{m}(\ell(x_{i},y_{i},s_{f_{1},h})-\ell(x_{i},y_{i},s_{f_{2},h}))^{2}$ (A.26)
@@
    $\sqrt{\frac{1}{m}\sum_{i=1}^{m}(\widetilde{\ell}(x_{i},y_{i},f_{1})-\widetilde{\ell}(x_{i},y_{i},f_{2}))^{2}}\leq 4M^{\frac{1}{2}}\|f_{1}-f_{2}\|_{L^{\infty}(\Omega_{R_{1}})}+8d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2).$ (A.28)
    For any $\varepsilon\geq 16d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$, according to A.3,
    $\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})},\varepsilon)$ $\leq\log\mathcal{N}(\mathcal{F},\|\cdot\|_{L^{\infty}(\Omega_{R_{1}})},\varepsilon/(8M^{\frac{1}{2}}))$ (A.29)
@@
    Plug in (A.25) and let $\theta=16d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$,
    $\mathcal{R}_{m}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}S_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)(R\vee T)M}{\varepsilon}\right)}{m}}\mathrm{d}\varepsilon\right)$ (A.30)
@@
    =:~$=:\widetilde{\mathcal{R}}_{m}(r)$
Combine the three steps above, by Lemma A.11 with $B_{0}=8M^{2}\exp(-C_{1}^{\prime}R^{2}),B=4M,b=M$, it holds that with probability no less than $2\displaystyle 1-2m\exp(-C_{1}^{\prime}R^{2})-\delta/$ 2, for any $f\in\mathcal{F}$,
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]$ $\leq\frac{2}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)+C_{5}M\left(r_{m}^{*}+\frac{\log(\log(m)/\delta)}{m}\right)$ (A.31)
@@
$\frac{1}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)$ $\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]+C_{5}M\left(r_{m}^{*}+\frac{\log(\log(m)/\delta)}{m}\right)$ (A.32)
@@
where $r_{m}^{*}$ is the largest fixed point of ~$\widetilde{\mathcal{R}}_{m}$, and it can be bounded as
$r_{m}^{*}\leq C_{6}\left(d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)+\frac{S_{f}L_{f}\log\left(mL_{f}W_{f}(B_{f}\vee 1)(R\vee T)M\right)}{m}\right),$ (A.33)
@@
$\left|\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]-\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]\right|\leq 2M\exp(-C_{1}^{\prime}R^{2}).$ (A.34)
Combine this with (A.31),(A.32),
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]$ (A.35)
@@
$\frac{1}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]$ $\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (A.36)
@@
Plug in the definition of $M=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T}+1\Big{)}\right)$ and let $R=C\log^{\frac{1}{2}}(md_{x}M_{f}/\delta)$ for some large constant $C$. Hence (A.35) and (A.36) reduce to
$\displaystyle\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]+C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(\log(m)/\delta)}{m}\right),$ (A.37)
$\displaystyle\frac{1}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]+C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(\log(m)/\delta)}{m}\right),$ (A.38)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f},h})]$ $\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{\widehat{f},h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]+2C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(1/\delta)}{m}\right)$ (A.39)
@@
Proposition A.5 (Prop. 3.3).
[…]
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}).$ (A.40)
@@
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}\ell^{\mathbb{P}^{k}}(x,y,s_{\widehat{f}^{k},\widehat{h}})\leq\inf_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{4}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}}(x,y,s_{f^{k},h})]+C_{Z}\log^{3}(nK/\delta)\left(r_{z}+\frac{\log(1/\delta)}{nK}\right),$ (A.41)
@@
$\Phi=\left\{z=(k,x,y)\mapsto\widetilde{\ell}(z,{\bm{f}},h):=(\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{k}))\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}:{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}\right\},$ (A.42)
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h).$ (A.43)
where $z_{i}^{k}=(k,x_{i}^{k},y_{i}^{k})$. Below we reason conditioned on this event and verify the conditions in Lemma A.12.
Following Step 1 and 2 in Proposition A.4, we have for any ${\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}$,
|~$|\widetilde{\ell}(z,{\bm{f}},h)|\leq M:=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big{(}\frac{\log(1/T_{0})}{T}+1\Big{)}\right).$ (A.44)
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\widetilde{\ell}(z^{k},{\bm{f}},h)^{2}]\leq\frac{4M}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\widetilde{\ell}(z^{k},{\bm{f}},h)]+8M^{2}\exp(-C_{1}^{\prime}R^{2}).$ (A.45)
@@
$\Big{\|}\frac{1}{\sqrt{nK}}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\frac{1}{\sqrt{nK}}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2})\Big{\|}_{\psi_{2}}\leq 4\|\widetilde{\ell}(\cdot,{\bm{f}}_{1},h_{1})-\widetilde{\ell}(\cdot,{\bm{f}}_{2},h_{2})\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},$ (A.46)
where ^$\widehat{\mathbb{P}}_{n}^{(K)}:=\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\delta_{z_{i}^{k}}$ and $\textbf{diam}\big{(}\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})}\big{)}\leq 2\sqrt{r}$. By Dudley’s bound [Van Handel, 2014; Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta>0$,
$\mathcal{R}_{K,n}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},\varepsilon)}{nK}}\ \mathrm{d}\varepsilon\right).$ (A.47)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2}))^{2}$ (A.48)

### v2

@@ B.2 Proofs of Transfer Learning
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f},h})]\leq 4\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]+C_{xy}\log^{3}(m)r_{x},$ (B.19)
@@
$\Phi=\{(x,y)\mapsto\widetilde{\ell}(x,y,f):=(\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}}))\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}:f\in\mathcal{F}\},$ (B.20)
where the truncation radius $R\geq 1$ will be defined later. It is easy to show that with probability no less than $1-2m\exp(-C_{1}^{\prime}R^{2})$, it holds that $\|x_{i}\|_{\infty}\leq R$ for all $1\leq i\leq m$. Hence by definition, the empirical minimizer also satisfies $\widehat{f}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\frac{1}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)$. Below we reason conditioned on this event and verify the conditions required in Lemma B.11.
@@
    $\widetilde{\ell}(x,y,f)\leq\mathbb{E}_{t,x_{t}|x}\|s_{f,h}(x_{t},y,t)-\nabla_{x}\log\phi_{t}(x_{t}|x)\|^{2}\lesssim M_{f}^{2}+d_{x}\Big(\frac{\log(1/T_{0})}{T-T_{0}}+1\Big).$ (B.21)
    And by Lemma B.10,
    $-\widetilde{\ell}(x,y,f)\leq\mathbb{E}_{t,x_{t}|x}\|s_{*}^{\mathbb{P}}(x_{t},y,t)-\nabla_{x}\log\phi_{t}(x_{t}|x)\|^{2}\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}\lesssim C_{X}^{{}^{\prime\prime}}R^{6}+d_{x}\Big(\frac{\log(1/T_{0})}{T}+1\Big).$ (B.22)
    Let $M:=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big(\frac{\log(1/T_{0})}{T}+1\Big)\right)$ and thus $|\widetilde{\ell}(x,y,f)|\leq M$.
@@
    $\mathbb{E}_{(x,y)\sim\mathbb{P}}\left[\mathbbm{1}_{\|x\|_{\infty}\leq R}\left(\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})\right)^{2}\right]$ (B.23)
@@
    $\Big\|\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\sigma_{i}\widetilde{\ell}(x_{i},y_{i},f_{1})-\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\sigma_{i}\widetilde{\ell}(x_{i},y_{i},f_{2})\Big\|_{\psi_{2}}\leq 4\|\widetilde{\ell}(\cdot,\cdot,f_{1})-\widetilde{\ell}(\cdot,\cdot,f_{2})\|_{L^{2}(\widehat{\mathbb{P}}_{m})},$ (B.24)
    where $\widehat{\mathbb{P}}_{m}:=\frac{1}{m}\sum_{i=1}^{m}\delta_{(x_{i},y_{i})}$. Define $\Phi_{r}:=\{\varphi\in\Phi:\frac{1}{m}\sum_{i=1}^{m}\varphi(x_{i},y_{i})^{2}\leq r\}$ and it is easy to show that $\textbf{diam}\big(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})}\big)\leq 2\sqrt{r}$. By Dudley’s bound [Van Handel, 2014, Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta>0$,
    $\mathcal{R}_{m}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})},\varepsilon)}{m}}\ \mathrm{d}\varepsilon\right).$ (B.25)
@@
    $\frac{1}{m}\sum_{i=1}^{m}(\widetilde{\ell}(x_{i},y_{i},f_{1})-\widetilde{\ell}(x_{i},y_{i},f_{2}))^{2}$ $=\frac{1}{m}\sum_{i=1}^{m}(\ell(x_{i},y_{i},s_{f_{1},h})-\ell(x_{i},y_{i},s_{f_{2},h}))^{2}$ (B.26)
@@
    $\sqrt{\frac{1}{m}\sum_{i=1}^{m}(\widetilde{\ell}(x_{i},y_{i},f_{1})-\widetilde{\ell}(x_{i},y_{i},f_{2}))^{2}}\leq 4M^{\frac{1}{2}}\|f_{1}-f_{2}\|_{L^{\infty}(\Omega_{R_{1}})}+8d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2).$ (B.28)
    For any $\varepsilon\geq 16d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$, according to B.3,
    $\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{m})},\varepsilon)$ $\leq\log\mathcal{N}(\mathcal{F},\|\cdot\|_{L^{\infty}(\Omega_{R_{1}})},\varepsilon/(8M^{\frac{1}{2}}))$ (B.29)
@@
    Plug in (B.25) and let $\theta=16d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)$,
    $\mathcal{R}_{m}(\Phi_{r})$ $\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{C_{4}S_{f}L_{f}\log\left(\frac{L_{f}W_{f}(B_{f}\vee 1)(R\vee T)M}{\varepsilon}\right)}{m}}\mathrm{d}\varepsilon\right)$ (B.30)
@@
    $=:\widetilde{\mathcal{R}}_{m}(r)$
Combine the three steps above, by Lemma B.11 with $B_{0}=8M^{2}\exp(-C_{1}^{\prime}R^{2}),B=4M,b=M$, it holds that with probability no less than $1-2m\exp(-C_{1}^{\prime}R^{2})-\delta/2$, for any $f\in\mathcal{F}$,
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]$ $\leq\frac{2}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)+C_{5}M\left(r_{m}^{*}+\frac{\log(\log(m)/\delta)}{m}\right)$ (B.31)
@@
$\frac{1}{m}\sum_{i=1}^{m}\widetilde{\ell}(x_{i},y_{i},f)$ $\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]+C_{5}M\left(r_{m}^{*}+\frac{\log(\log(m)/\delta)}{m}\right)$ (B.32)
@@
where $r_{m}^{*}$ is the largest fixed point of $\widetilde{\mathcal{R}}_{m}$, and it can be bounded as
$r_{m}^{*}\leq C_{6}\left(d_{x}^{\frac{1}{2}}M\exp(-C_{0}^{\prime}R^{2}/2)+\frac{S_{f}L_{f}\log\left(mL_{f}W_{f}(B_{f}\vee 1)(R\vee T)M\right)}{m}\right),$ (B.33)
@@
$\left|\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]-\mathbb{E}_{(x,y)\sim\mathbb{P}}[\widetilde{\ell}(x,y,f)]\right|\leq 2M\exp(-C_{1}^{\prime}R^{2}).$ (B.34)
Combine this with (B.31),(B.32),
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ $\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]$ (B.35)
@@
$\frac{1}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]$ $\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell(x,y,s_{f,h})-\ell(x,y,s_{*}^{\mathbb{P}})]$ (B.36)
@@
Plug in the definition of $M=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big(\frac{\log(1/T_{0})}{T}+1\Big)\right)$ and let $R=C\log^{\frac{1}{2}}(md_{x}M_{f}/\delta)$ for some large constant $C$. Hence (B.35) and (B.36) reduce to
$\displaystyle\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]+C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(\log(m)/\delta)}{m}\right),$ (B.37)
$\displaystyle\frac{1}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{f,h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]\leq 2\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]+C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(\log(m)/\delta)}{m}\right),$ (B.38)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f},h})]$ $\leq\frac{2}{m}\sum_{i=1}^{m}[\ell(x_{i},y_{i},s_{\widehat{f},h})-\ell(x_{i},y_{i},s_{*}^{\mathbb{P}})]+2C_{7}M_{f}^{2}\log^{3}(m/\delta)\left(r_{m}^{\dagger}+\frac{\log(1/\delta)}{m}\right)$ (B.39)
@@
Proposition B.5 (Prop. 3.3).
There exists some constant $C_{Z},C_{R}$ such that the following holds. For any $\mathbb{P}^{1},\cdots,\mathbb{P}^{K}$, let $x_{1}^{k},\cdots,x_{n}^{k}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}^{k}$ for any $k$ and $(x_{i}^{k})_{i,k}$ are all independent. Consider the empirical minimizer
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\ell(x_{i}^{k},y_{i}^{k},s_{f^{k},h}).$ (B.40)
@@
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}\ell^{\mathbb{P}^{k}}(x,y,s_{\widehat{f}^{k},\widehat{h}})\leq\inf_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{4}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}}(x,y,s_{f^{k},h})]+C_{Z}\log^{3}(nK/\delta)\left(r_{z}+\frac{\log(1/\delta)}{nK}\right),$ (B.41)
@@
$\Phi=\left\{z=(k,x,y)\mapsto\widetilde{\ell}(z,{\bm{f}},h):=(\ell(x,y,s_{f^{k},h})-\ell(x,y,s_{*}^{k}))\cdot\mathbbm{1}_{\|x\|_{\infty}\leq R}:{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}\right\},$ (B.42)
@@
$\widehat{{\bm{f}}},\widehat{h}=\operatorname*{arg\,min}_{{\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\widetilde{\ell}(z_{i}^{k},{\bm{f}},h).$ (B.43)
where $z_{i}^{k}=(k,x_{i}^{k},y_{i}^{k})$. Below we reason conditioned on this event and verify the conditions in Lemma B.12.
Following Step 1 and 2 in Proposition B.4, we have for any ${\bm{f}}\in\mathcal{F}^{\otimes K},h\in\mathcal{H}$,
$|\widetilde{\ell}(z,{\bm{f}},h)|\leq M:=C\left(C_{X}^{{}^{\prime\prime}}R^{6}+M_{f}^{2}+d_{x}\Big(\frac{\log(1/T_{0})}{T}+1\Big)\right).$ (B.44)
$\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\widetilde{\ell}(z^{k},{\bm{f}},h)^{2}]\leq\frac{4M}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\widetilde{\ell}(z^{k},{\bm{f}},h)]+8M^{2}\exp(-C_{1}^{\prime}R^{2}).$ (B.45)
@@
$\Big\|\frac{1}{\sqrt{nK}}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\frac{1}{\sqrt{nK}}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2})\Big\|_{\psi_{2}}\leq 4\|\widetilde{\ell}(\cdot,{\bm{f}}_{1},h_{1})-\widetilde{\ell}(\cdot,{\bm{f}}_{2},h_{2})\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},$ (B.46)
where $\widehat{\mathbb{P}}_{n}^{(K)}:=\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\delta_{z_{i}^{k}}$ and $\textbf{diam}\big(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})}\big)\leq 2\sqrt{r}$. By Dudley’s bound [Van Handel, 2014, Wainwright, 2019], there exists an absolute constant $C_{0}$ such that for any $\theta>0$,
$\mathcal{R}_{K,n}(\Phi_{r})\leq C_{0}\left(\theta+\int_{\theta}^{2\sqrt{r}}\sqrt{\frac{\log\mathcal{N}(\Phi_{r},\|\cdot\|_{L^{2}(\widehat{\mathbb{P}}_{n}^{(K)})},\varepsilon)}{nK}}\ \mathrm{d}\varepsilon\right).$ (B.47)
@@
$\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{1},h_{1})-\widetilde{\ell}(z_{i}^{k},{\bm{f}}_{2},h_{2}))^{2}$ (B.48)